# Brain functions emerge as thermal equilibrium states of the connectome

### v2

@@ Results
10. Asymmetrical integration. Some neuron classes show asymmetry in their integration capacity. For instance, the touch receptor neuron ALMR is in group 1 while its sister cell ALML belongs to group 2 (see Tab. 3). The polymodal amphid sensory neuron class ASH, which has been shown [31, 22, 26] to have a multisensory integration function, has its right cell ASHR in group 1 and left one ASHL in group 2. While PHAL and PHAR are both in group 3, PHBL belongs to group 1 and its sister cell PHBR is in group 3. This indicates that among the phasmid neuron classes PHA and PHB, which have been shown [26, 28] to belong to an antagonistic functional circuit that integrates sensory responses from the amphids ASH and ASK to mediate chemo-repulsion behavior, PHBR is structurally well positioned to have greater long-range integration function.
@@
11. IC predicts response of AIY in AFD– or AWC–ablated animals. As a direct application of the $\operatorname{\bf IC}$ measure, we theoretically reproduced an experimental study by Kano et al. [30] that has shown that the thermosensory neuron class AWC regulates the information processing in the AFD-AWC-AIY circuit [35, 34], which is part of the functional circuit mediating the C. elegans’thermotaxis’ behavior [39, 37].
@@ Discussion
From the interconnectivity matrices $[\operatorname{{\bf x}}^{\bullet|\beta}]$ formed by the pure functional states, we extracted the directed weighted networks $\mathcal{F}^{\beta}$—the pure functional connectomes—consisting of all the neural emittances that are PFCs; i.e., determined by the network topology of the structural connectome. In particular, at a functional inverse temperature $\beta_{f}=4.51$, $\mathcal{F}^{\beta}$ represents only about $12\%$ of all possible neural emittances. Extrapolating this result to undirected graphs, this connectivity density would account for approximately $24\%$ of functional connections, which is surprisingly consistent with functional connectivity thresholding in neuroimaging studies in humans and animals published in the literature [60].
@@
Promising engineering techniques have recently been developed for emulating synaptic plasticity and exploring the structure-function relationships [46, 48]. These observational approaches involve inserting synthetic synapses into in vivo neural circuits and investigating how the added connections interact with brain functional properties and modify animal behavior. However, these synaptic manipulations are performed under plasticity hypotheses linking local structural modifications to changes in global information flow and behavioral adaptation [52]. Moreover, while (3) shows that adding a direct connection from neuron $u$ to $v$ strengthens the neural emittance of $u$ to $v$, we observed that structural connections do not necessarily underlie PFCs, and conversely, PFCs do not always align with structural connections. This implies that not all insertions of synthetic synapses are likely to generate functional connections. Therefore, although synaptic plasticity was not the focus of this work, our framework can guide the identification of changes in neuronal structural interconnectivities that lead to functional network reconfigurations, offering a practical atlas of relevant neuron pairs for more efficient synaptic engineering.
@@
On the other hand, experimental research in C. elegans has identified functional interactions among neurons that differ from predictions based on anatomical information flow pathways, with evidence showing that extrasynaptic transmission contributes to these differences [50]. This indicates that direct connections via extracellular neurotransmitters [8] dynamically and selectively add to the original graph, altering its computational and functional properties by establishing new functional connections. This property can be simulated by extracting from the connectivity matrix $[\operatorname{{\bf x}}^{\bullet|\beta}]$ neural emittance connections that (i) are not mapped to existing structural connections and (ii) do not belong to the pure functional connectome $\mathcal{F}^{\beta}$. New direct edges representing virtual extrasynaptic connections are then added between the corresponding neuron pairs to investigate their impact on the pure functional connectome. Indeed, performing such a simulation on RID revealed the emergence of new PFCs that correspond with the extrasynaptic signaling-based functional connectivity identified by Randi et al. [50].
@@
Methods
The C. elegans somatic connectome. We have merged two datasets of the somatic connectome of the hermaphrodite worm C. elegans that are publicly available on the WormAtlas [2]: the first dataset, Data1, is from Varshney et al. [61], and the second one, Data2, is from the recently published serial electron microscopy reconstructions by Cook et al. [13]. Specifically, we have kept all synaptic connections from Data1 (279 connected neurons and 8171 chemical and electrical synapses), to which we have added all synaptic connections in Data2 that were not originally in Data1 (3900 additional connections). The resulting connectome consists of 280 connected neurons and 12071 synaptic connections. Note that in the old version (Data1), VC06 had no connections, while it is connected in the revised connectome.
Computing the structure-function divergence and the mean total receptance. To calculate $\operatorname{\bf sfd}(v,\beta)$ for a neuron from Eq. 5, we excluded self-loops by using the connectivity distribution $\bar{\operatorname{{\bf k}}}^{v}$ instead of $\operatorname{{\bf k}}^{v}$, where $\bar{\operatorname{{\bf k}}}^{v}_{v}=0$ and for $u\neq v$, $\bar{\operatorname{{\bf k}}}^{v}_{u}=\operatorname{{\bf k}}^{v}_{u}/\sum_{w\neq v}\operatorname{{\bf k}}^{v}_{w}$ if $\operatorname{{\bf k}}^{v}_{u}>0$.
Statistical significance of neural emittance. In order to know the extent to which a given neural $\beta$–emittance is statistically significant, we used a bootstrapping technique consisting of generating 5000 random directed graphs with the same degree sequence as the graph $G$ representing the connectome, and computing the neural emittance of the same neuron in each generated graph at the same inverse temperature $\beta$. The neural emittance is then statistically significant if it occurs with the same or larger value in less than $5\%$ of the time ($p<0.05$).
Data and code availability

### v3

@@ III.4 Classifying neurons from IC
Table 2: Classification of neurons by IC
III.5 Asymmetrical integration
Some neuron classes show asymmetry in their integration capacity. For instance, the touch receptor neuron ALMR is in group 1 while its sister cell ALML belongs to group 2 (see Tab. 2). The polymodal amphid sensory neuron class ASH, which has been shown [35, 58, 59] to have a multisensory integration function, has its right cell ASHR in group 1 and left one ASHL in group 2. While PHAL and PHAR are both in group 3, PHBL belongs to group 1 and its sister cell PHBR is in group 3. This indicates that among the phasmid neuron classes PHA and PHB, which have been shown [59, 60] to belong to an antagonistic functional circuit that integrates sensory responses from the amphids ASH and ASK to mediate chemo-repulsion behavior, PHBR is structurally well positioned to have greater long-range integration function.
IV Experimental validation
IV.1 PTFC reveal entrality of the locomotory circuit
We investigated $\mathcal{F}^{\beta}$ at the functional $\beta_{f}=4.51$ (see Supplementary Data 3 for the edgelist [39]) and found that it reveals connectivity patterns that coincide with circuitries that have been reported by researchers to be the functional substrates of well studied C. elegans complex behaviors such as locomotion and touch-induced movement [13, 36], indicating that our network model provides a theoretical basis for understanding the functional organization of a nervous system.
Indeed, $\mathcal{F}^{\beta}$ has 8932 weighted directed edges, compared to the network $\mathcal{K}$ formed by the structural connectivity of all neurons which has 4927 weighted directed edges (see Supplementary Data 2 [39]). It represents $11.68\%$ of all the potential functional connections given by the adjacency matrix $\operatorname{{\bf x}}^{\bullet|\beta}]$, implying that only about $12\%$ of all neural $\beta$–emittance connections are PTFCs. This network is schematically represented in Figure 3.
We observed a substantial topological discordance between $\mathcal{F}^{\beta}$ and $\mathcal{K}$. For instance, the motor neuron classes AS, VA, DA, VB, and DB, have the highest in-degree centrality $\mathcal{F}$, with AS08, DA06, AS07, DA07, VA08, DB05, VA10, DB06, AS09, VA11 being the 10 most central neurons in terms of incoming connections, and specifically AS08, which is ranked 278 in the anatomical network (see Table 3), receiving 219 PTFCs. Notably, none of these neurons is among the most in-connected in the structural connectome ($k^{in}=2$ for AS08 and 7 for DA06) which has the command interneuron classes AVA, AVB, AVE, AVD, and PVC receiving the highest numbers of incoming connections. Moreover, VA08 has the highest weighted in-degree $w^{in}$ followed by DA06, and DB05 has the largest weighted out-degree $w^{out}$, while they are poorly connected anatomically. Additionally, some of these neurons are highly ranked in terms of out-degree and weighted out-degree (see Table 3). Specifically, DA07 has the second largest number of outgoing connections ($k^{ou}=64$), compared to its $3$ anatomical weighted connections, VB07 has out-degree rank 6 and weighted out-degree rank 7, and DB05 has out-degree rank 9 and weighted out-degree rank 1. Interestingly, PVCL and PVCR are the most functionally in- and out-connected among the command interneurons most of whom have their anatomical status downgraded in the pure functional connectome. Indeed, on the one hand, AVAL, AVAR, PVCL, and PVCR have their respective structural weighted in-degree ranks almost maintained in $\mathcal{F}^{\beta}$, while their weighted out-degree ranks are not, and on the other, PVCL and PVCR have their high structural out-degree ranks almost maintained in $\mathcal{F}^{\beta}$ while their weighted out-degree ranks are not.
IV.2 Mechanoreceptor neurons have higher functional out-degree
Additionally, we found that the posterior mechanoreceptor neuron [54] PLML is the most central in number of outgoing connections in $\mathcal{F}^{\beta}$ ($k^{out}=66$), and the anterior mechanoreceptor ALMR is among the 30 most out-connected with $k^{out}=54$ in $\mathcal{F}^{\beta}$ (ranked 27 in Table 3). By comparison, both neurons have very low out-degrees of respectively 4 and 12 and low ranks of 272 and 183 in the structural network $\mathcal{K}$, and both neurons have low ranks in terms of weighted out-degree (259 and 247, respectively). More generally, many mechanoreceptor neurons [35, 54, 62], including ALA, ASHR, AVM, DVA, FLPL/R, OLQVL, etc., are among the most out-connected in $\mathcal{F}^{\beta}$.
@@ IV.2 Mechanoreceptor neurons have higher functional out-degree
IV.3 IC predicts response of AIY in AFD– or AWC–ablated animals
As a direct application of the $\operatorname{\bf IC}$ measure, we theoretically reproduced an experimental study by Kano et al. [63] that has shown that the thermosensory neuron class AWC regulates the information processing in the AFD-AWC-AIY circuit [56, 34], which is part of the functional circuit mediating the C. elegans’thermotaxis’ behavior [33, 57].
@@ V Discussion
From the interconnectivity matrices $\operatorname{{\bf x}}^{\bullet|\beta}]$ formed by the pure functional states, we extracted the directed weighted networks $\mathcal{F}^{\beta}$—the pure functional connectomes—consisting of all the neural emittances that are PTFCs; i.e., determined by the network topology of the structural connectome. In particular, at a functional inverse temperature $\beta_{f}=4.51$, $\mathcal{F}^{\beta}$ represents only about $12\%$ of all possible neural emittances. Extrapolating this result to undirected graphs, this connectivity density would account for approximately $24\%$ of functional connections, which is surprisingly consistent with functional connectivity thresholding in neuroimaging studies in humans and animals published in the literature [64].
@@
Promising engineering techniques have recently been developed for emulating synaptic plasticity and exploring the structure-function relationships [65, 66]. These observational approaches involve inserting synthetic synapses into in vivo neural circuits and investigating how the added connections interact with brain functional properties and modify animal behavior. However, these synaptic manipulations are performed under plasticity hypotheses linking local structural modifications to changes in global information flow and behavioral adaptation [67]. Moreover, while (3) shows that adding a direct connection from neuron $u$ to $v$ strengthens the neural emittance of $u$ to $v$, we observed that structural connections do not necessarily underlie PTFCs, and conversely, PTFCs do not always align with structural connections. This implies that not all insertions of synthetic synapses are likely to generate functional connections. Therefore, although synaptic plasticity was not the focus of this work, our framework can guide the identification of changes in neuronal structural interconnectivities that lead to functional network reconfigurations, offering a practical atlas of relevant neuron pairs for more efficient synaptic engineering.
@@
On the other hand, experimental research in C. elegans has identified functional interactions among neurons that differ from predictions based on anatomical information flow pathways, with evidence showing that extrasynaptic transmission contributes to these differences [52]. This indicates that direct connections via extracellular neurotransmitters [50] dynamically and selectively add to the original graph, altering its computational and functional properties by establishing new functional connections. This property can be simulated by extracting from the connectivity matrix $\operatorname{{\bf x}}^{\bullet|\beta}]$ neural emittance connections that (i) are not mapped to existing structural connections and (ii) do not belong to the pure functional connectome $\mathcal{F}^{\beta}$. New direct edges representing virtual extrasynaptic connections are then added between the corresponding neuron pairs to investigate their impact on the pure functional connectome. Indeed, performing such a simulation on RID revealed the emergence of new PTFCs that correspond with the extrasynaptic signaling-based functional connectivity identified by Randi et al. [52].
@@
VI Materials and Methods
The C. elegans somatic connectome.
We have merged two datasets of the somatic connectome of the hermaphrodite worm C. elegans that are publicly available on the WormAtlas [38]: the first dataset, Data1, is from Varshney et al. [26], and the second one, Data2, is from the recently published serial electron microscopy reconstructions by Cook et al. [14]. Specifically, we have kept all synaptic connections from Data1 (279 connected neurons and 8171 chemical and electrical synapses), to which we have added all synaptic connections in Data2 that were not originally in Data1 (3900 additional connections). The resulting connectome consists of 280 connected neurons and 12071 synaptic connections. Note that in the old version (Data1), VC06 had no connections, while it is connected in the revised connectome.
Computing the structure-function divergence and the mean total receptance.
To calculate $\operatorname{\bf sfd}(v,\beta)$ for a neuron from Eq. 5, we excluded self-loops by using the connectivity distribution $\bar{\operatorname{{\bf k}}}^{v}$ instead of $\operatorname{{\bf k}}^{v}$, where $\bar{\operatorname{{\bf k}}}^{v}_{v}=0$ and for $u\neq v$, $\bar{\operatorname{{\bf k}}}^{v}_{u}=\operatorname{{\bf k}}^{v}_{u}/\sum_{w\neq v}\operatorname{{\bf k}}^{v}_{w}$ if $\operatorname{{\bf k}}^{v}_{u}>0$.
Statistical significance of neural emittance.
In order to know the extent to which a given neural $\beta$–emittance is statistically significant, we used a bootstrapping technique consisting of generating 5000 random directed graphs with the same degree sequence as the graph $G$ representing the connectome, and computing the neural emittance of the same neuron in each generated graph at the same inverse temperature $\beta$. The neural emittance is then statistically significant if it occurs with the same or larger value in less than $5\%$ of the time ($p<0.05$).
Data and code availability.
@@ VI Materials and Methods
Acknowledgment
This work was supported by the Natural Sciences and Engineering Research Council of Canada through the CRC grant NC0981.